OpenStudy (anonymous):

Alright, I need somebody to walk me through this. Rewrite the equation in vertex form. Name the vertex and y-intercept. y = x2 – 10x + 15 I VAGUELY understand the concept... But I could really use some direction.

OpenStudy (anonymous):

Do you know how to complete the square?

OpenStudy (anonymous):

I think so... one second...

OpenStudy (anonymous):

okay, I have (x=5)^2=10. Is that right so far?

OpenStudy (anonymous):

Is it (x-5)^2=10?

OpenStudy (anonymous):

oh I meant a + but could it be either one?

OpenStudy (anonymous):

It should be a minus... y=x^2-10x+25-25+15 y=(x-5)^2-25+15 y=(x-5)^2-10

OpenStudy (anonymous):

\[x + 5 =\sqrt{10}\] is that correct?

OpenStudy (anonymous):

No you stop when you have the vertex form y=a(x-h)^2+k

OpenStudy (anonymous):

Idk what vertex form is. :P okay so... y=(x-5)^2-10 is the final vertex form answer?

OpenStudy (anonymous):

Yes that is vertex form. its in the form y=a(x-h)^2+k where (h,k) is the vertex so what would the vertex be?

OpenStudy (anonymous):

x=5 y=-10

OpenStudy (anonymous):

So vertex is (5,-10) And to find the y intercept you just input 0 into the equation for x and solve for y

OpenStudy (anonymous):

so y-intercept would be (0,15) ? :D

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Awesome! That is not as hard as I was making it! Thank you!

OpenStudy (anonymous):

You're welcome :)

OpenStudy (anonymous):

@CalebBeavers What do I do if there is a fraction in front of my x^2?

OpenStudy (jojomojo):